# Learned Hemodynamic Coupling Inference in Resting-State Functional MRI

### v1

@@ III-D2 Calibration of Simulator
where $(\theta_{k},s_{k},\epsilon_{k})\sim p(\theta,s,\epsilon|\nu)$, for k=1,,Kk=1,...,K. Clearly, (19) is not differentiable in $\nu$, so we use a zero-th order optimizer to estimate the minimizer [32].
III-E Uncertainty Quantification
@@
We therefore propose a double bootstrap based approach, adapting the method from [33], originally developed for nonparametric regression, to our current setting. The algorithm is outlined below.
@@
    Generate $B$ resamplings of the observed signals $Y^{(b)}$, b=1,,Bb=1,...,B using the stationary bootstrap [34]. Unlike the classical bootstrap for independent data, this method resamples temporally contiguous blocks of observations to approximately preserve the time dependence structure of the signals.
@@
IV Experiments and Implementation
IV-A Hemodynamic Models
Figure 1: Hemodynamic response kernel  (20) for several $\theta$.
@@ IV-A Hemodynamic Models
with constants $a_{1}=6$, $a_{2}=16$, $b_{1}=1$, $b_{2}=1$, $c=1/6$. The first model is a one-parameter shifted double gamma model [18], given by
@@
with parameter bounds $\theta\in[0.5,2.5]$. Figure 1 shows how the kernel shape changes with respect to $\theta$. The second model is a two-parameter basis expansion of the form
@@
for which reasonable physiological ranges are $\theta_{1}\in[0.2,2.0]$, $\theta_{2}\in[-1.0,1.0]$ [5].
IV-B Datasets
IV-B1 In-vivo Data
@@
We use data from the publicly available Young Adult Human Connectome Project (HCP-YA) [35] to evaluate our method. Specifically, we use the preprocessed rsfMRI data mapped to the left cortical surface mesh with $32,492$ vertices and registered using FreeSurfer. Details on the acquisition and preprocessing protocols can be found in the standard reference [36], and are thus omitted here.
IV-B2 Synthetic Data
@@
A synthetic hemodynamic parameter field $\tilde{\theta}$ is sampled from the GP prior outlined in Section III-B, with $\kappa=5\cdot 10^{-3}$, $\tau^{2}=10^{4}$ for the 1-parameter model (20) and $\kappa_{1}=\kappa_{2}=5\cdot 10^{-2}$, $\tau_{1}^{2}=\tau^{2}_{2}=10^{4}$ for the 2-parameter model (21). The calibration procedure from Section III-D2 was run on the rsfMRI data of a randomly selected HCP-YA subject under both hemodynamic forward models, and the results were used to parameterize the neural signal simulator (Section II-A) and measurement error model (Section II-C). Using the calibrated simulator and hemodynamic parameter field, we generated synthetic rsfMRI time series at all $V=32,492$ left cortical surface mesh vertices for $M=1200$ time points with $t_{r}=0.72$s, matching the HCP-YA protocol.
IV-C Implementation details
@@
For the summary network $T_{\psi}(\cdot)$, we use a fully connected network with ReLU activations, with an initial non-trainable encoding layer which takes the Fourier transform of the BOLD signals $y$ at the set of Nyquist frequencies. We set the number of hidden layers to be 3, with hidden widths $M,\lfloor\frac{M}{2}\rfloor,\lfloor\frac{M}{4}\rfloor$. For the marginal likelihood emulator $p_{\gamma}(\cdot|\cdot)$, we use a conditional neural spline normalizing flow [37] with $5$ spline coupling transforms, each with a fully connected 3 layer conditioning network with hidden widths of 64 and ReLU activations. Both networks are trained on synthetic data generated from the simulator outlined in Section II using Adam with a batch size of $100$ and learning rate of $10^{-5}$ for $10^{5}$ iterations. Note that separate summary network and likelihood model pairs were trained for the forward models (20) and (21).
@@
For calibrating the simulator hyperparameters using the non-differentiable loss (19), we used Bayesian optimization [32] under the expected improvement acquisition function for a total of $200$ runs using the rsfMRI from a randomly selected HCP-YA subject and both forward models. The prior smoothness hyperparameters are selected by maximizing (17) over a pre-defined grid. For single subject analysis, this was repeated for each subject. For multi-subject analysis, we use a common set of smoothing parameters for all reconstructions, as is standard practice.
IV-D Competing Hemodynamic Estimators
@@
where $A(\theta)\in\mathbb{R}^{M\times M}$ is a convolution matrix formed from the kernel (20), $s\in\mathbb{R}^{M}$ denotes the discretized neural signals, and $\eta>0$ is a regularization parameter. Optimization is performed using block-coordinate descent, where the (convex) subproblem in $s$ is solved using alternating direction method of multipliers and the (non-convex) subproblem in $\theta$ is solved using golden section search. We set the penalty parameter $\eta=\sigma\sqrt{2\log(M)}$, where $\sigma$ is the calibrated noise level. The second approach, referred to as DUP (deep unrolled prior), adapts [26] to perform vertex-wise joint reconstruction. Specifically, DUP first pre-trains an unrolled network $\mathcal{A}_{\rho}:\mathbb{R}^{M}\times\mathbb{R}^{M\times M}\mapsto\mathbb{R}^{M}$, parameterized by weights $\rho$, to minimize
@@
via golden section search. We take $\mathcal{A}_{\rho}$ to be a 5-layer unrolled fully connected ReLU network with hidden width $M$ and train on synthetic data from the simulator model in Section II using Adam (learning rate $10^{-5}$, $10^{5}$ iterations, batch size $100$).
For the two-parameter model (21), we compare our method to [19], implemented in the MATLAB package rsHRF [38]. rsHRF uses a two stage procedure, first estimating neural spike times by thresholding the fMRI time series, then estimating hemodynamic parameters conditional on the detected spikes. We used the suggested defaults for all method hyperparameters.
TABLE I: Synthetic-data results: estimation performance (MSE, bias) and uncertainty quantification (empirical coverage (EC) and average interval length), averaged across the cortical surface.
@@
IV-E Evaluation
IV-E1 Hemodynamic Reconstruction
To evaluate the performance of the hemodynamic estimation on synthetic data, we compute the mean squared error (MSE) and bias of the reconstructions, averaged over the mesh. The quality of the uncertainty quantification procedure from Section III-E is assessed by the empirical coverage of the $95\%$ pointwise intervals. In the synthetic data experiments, this is computed as the fraction of mesh locations for which the intervals cover the true hemodynamic parameter. For in-vivo data, where we lack access to ground truth, we instead approximate the empirical coverage of the intervals using the HCP test-retest data; where one scan is used to form the intervals and their empirical coverage is calculated over the reconstructions from the other scan. Because hemodynamic response shape is expected to be relatively stable within healthy subjects between multiple scans over a short time period (months), the test-retest scans can be (approximately) considered as independent replications from model (4) with the same $\theta$, though different $s$ and $\epsilon$, and thus can be used to assess the stability of our hemodynamic coupling estimation methodology. We also compute the average interval length over the mesh for use as a scalar measure of reconstruction uncertainty.
IV-E2 Connectivity
Spatial variability in the cortical hemodynamic response can bias downstream functional connectivity estimates [5, 6]. We therefore evaluate whether our spatially varying hemodynamic estimates can improve connectivity analysis by comparing seed-based effective connectivity computed from three signal representations: I) raw (non-deconvolved) BOLD signals, II) deconvolved BOLD signals assuming a spatially constant and known hemodynamic response, III) deconvolved BOLD signals under a spatially varying hemodynamic response estimated by our methodology. Under the LTI forward model (3), with a known response kernel (either fixed or pre-estimated), the resulting linear inverse problem for latent signal reconstruction has been widely studied. For simplicity, we choose to perform deconvolution using the Wiener filter [39], acknowledging that better performance may be obtained through more modern techniques. Using a predefined seed region of interest (ROI), we estimate the effective connectivity from the time series at each non-ROI vertex to the ROI-averaged time series. We use a bivariate Granger Causality (GC) based definition of effective connectivity [40] with an autoregressive lag of $2$. For each non-ROI vertex, p-values are calculated from the $F$-tests assessing GC and then corrected to control false discovery rate at the $0.05$ level using the Benjamini-Hochberg procedure [41].
IV-E3 Population Analysis
We evaluate inter-subject variation of the hemodynamic field estimates using the functional principal components analysis (fPCA) approach proposed in [42]. For a cohort of $N$ subjects, fPCA is used to obtain the Karhunen–Loève decomposition of the scalar hemodynamic coupling parameter $\tilde{\theta}_{i}(x_{v})$ for subject $i$ at location $x_{v}$:
V Results
V-A Synthetic Data Results
Figure 2: Estimate of the $\theta$-field for the one-parameter hemodynamic coupling model (20) on the HCP test-retest data for three randomly selected subjects. Inter-subject variability is noticeably larger than within-subject between-scan variability.
@@ V-A Synthetic Data Results
Table I A) compares the estimation results for the synthetic data using the one-parameter model (20). Our method exhibits mean squared errors and biases that are, on average, an order of magnitude smaller than both JointMAP and DUP. Recall that both of the competitors perform joint reconstruction of signal and hemodynamic parameters, resulting in the estimation of the latter being dependent on the estimation of the former. Since even the non-blind signal reconstruction is ill-posed in this setting, we attribute the poorer performance of the competing methods, at least in part, to the propagation of errors from the signal reconstruction to the target hemodynamic estimation. Our approach avoids this issue by marginalizing the signal out of the data likelihood.
@@
The second key differentiating feature of our method is the incorporation of spatial regularization. To evaluate its effect, we drop the prior from Section III-B and apply the learned marginal posterior mean (6), denoted MPM, independently to each mesh BOLD time series. As shown in the third row of Table I A), on average, this results in a greater than $6\times$ increase in the MSE compared to our spatially regularized framework. While MPM outperforms both JointMAP and DUP owing to its avoidance of explicit signal reconstruction, we see that incorporating spatial smoothness of the hemodynamics provides substantial additional improvement.
@@
Table I B) shows the estimation results for the two-parameter forward model (21). Estimation of $\theta_{1}$ echoes the results from Table I A), where our method again results in an order of magnitude reduction in average MSE compared to rsHRF. Although our method exhibits lower MSE for $\theta_{2}$ as well, both approaches perform relatively poorly for this parameter.
@@
Table I C) shows the $95\%$ empirical pointwise coverage and average interval length for the procedure from Section III-E for both synthetic datasets. For the one-parameter model (20), we see reasonably good empirical coverage, with slight overconfidence. For the two-parameter model (21), the coverage on $\theta_{1}$ indicates a bit of conservatism, though the intervals are quite tight, indicating strong statistical identifiability of this parameter. While the empirical coverage for $\theta_{2}$ is reasonable, the estimation uncertainty is excessively large. Indeed, given $\theta_{2}\in[-1,1]$, the average interval length occupies approximately $70\%$ of the possible values of the parameter. Coupled with the relatively poor estimation performance for both methods for $\theta_{2}$, these results imply there are likely fundamental limitations on its statistical identifiability in resting-state data under HCP-style acquisition and noise level.
V-B In-vivo Data Results
@@
In this section, we apply our methodology under the one-parameter model (20) to the real resting-state HCP data. We choose this model over model (21) due to the superior statistical identifiability of model parameters reported in Section V-A.
V-B1 Hemodynamic Reproducibility
@@
The spatial maps of $\widehat{\boldsymbol{\theta}}$ for 3 randomly selected HCP test-retest subjects are shown in Figure 2. Qualitatively, we see that between-subject variability is noticeably larger than within-subject (between-scan) variability for each subject, with many unique subject-specific spatial features preserved across scans. This result is also apparent quantitatively, as the average within-subject squared norm of the difference between scan 1 and scan 2 ($\approx 0.0214$) was substantially lower than the average between-subject squared norm ($\approx 0.0590$).
@@
For each subject in Figure 2, we ran the uncertainty quantification procedure from Section III-E on the scan 1 data to construct pointwise $95\%$ intervals for the reconstructed field. We evaluated the empirical coverage of these intervals on the estimates from the scan 2 data. The average empirical coverage was $\approx 98.5\%$, indicating good calibration with mild conservatism.
V-B2 Population Analysis

### v2

@@ 3.4.2 Calibration of Simulator
where $(\theta_{k},s_{k},\epsilon_{k})\sim p(\theta,s,\epsilon|\nu)$, for k=1,,Kk=1,...,K. Clearly, (19) is not differentiable in $\nu$, so we use a zero-th order optimizer to estimate the minimizer (Frazier, 2018).
3.5 Uncertainty Quantification
@@
We therefore propose a double bootstrap based approach, adapting the method from Hall and Horowitz (2013), originally developed for nonparametric regression, to our current setting. The algorithm is outlined below.
@@
    Generate $B$ resamplings of the observed signals $Y^{(b)}$, b=1,,Bb=1,...,B using the stationary bootstrap (Politis and Romano, 1994). Unlike the classical bootstrap for independent data, this method resamples temporally contiguous blocks of observations to approximately preserve the time dependence structure of the signals.
@@
4 Experiments and Implementation
4.1 Hemodynamic Models
Figure 2: Hemodynamic response kernel for several $\theta$ for model (20) (left panel) and (21) (right panel).
@@ 4.1 Hemodynamic Models
with constants $a_{1}=6$, $a_{2}=16$, $b_{1}=1$, $b_{2}=1$, $c=1/6$. The first model is a one-parameter shifted double gamma model (Cherkaoui et al., 2021), given by
@@
with parameter bounds $\theta\in[0.5,2.5]$. The second model is a two-parameter basis expansion of the form
@@
for which reasonable physiological ranges are $\theta_{1}\in[0.2,2.0]$, $\theta_{2}\in[-1.0,1.0]$ (Lindquist et al., 2009). Figure 2 shows how the kernel shape changes with respect to $\theta$ for both models.
4.2 Datasets
4.2.1 In-vivo Data
@@
We use data from the publicly available Young Adult Human Connectome Project (HCP-YA) (Glasser et al., 2016) to evaluate our method. Specifically, we use the preprocessed rsfMRI data, denoised with ICA-FIX (Salimi-Khorshidi et al., 2014), and mapped to the left cortical surface mesh with $32,492$ vertices and registered using FreeSurfer. Details on the acquisition and preprocessing protocols can be found in the standard reference (Glasser et al., 2013), and are thus omitted here.
4.2.2 Synthetic Data
@@
A synthetic hemodynamic parameter field $\tilde{\theta}$ is sampled from the GP prior outlined in Section 3.2, with $\kappa=5\cdot 10^{-3}$, $\tau^{2}=10^{4}$ for the 1-parameter model (20) and $\kappa_{1}=\kappa_{2}=5\cdot 10^{-2}$, $\tau_{1}^{2}=\tau^{2}_{2}=10^{4}$ for the 2-parameter model (21). The calibration procedure from Section 3.4.2 was run on the rsfMRI data of a randomly selected HCP-YA subject under both hemodynamic forward models, and the results were used to parameterize the neural signal simulator (Section 2.1) and measurement error model (Section 2.3). Using the calibrated simulator and hemodynamic parameter field, we generated synthetic rsfMRI time series at all $V=32,492$ left cortical surface mesh vertices for $M=1200$ time points with $t_{r}=0.72$s, matching the HCP-YA protocol.
4.3 Implementation details
@@
For the summary network $T_{\psi}(\cdot)$, we use a fully connected network with ReLU activations, with an initial non-trainable encoding layer which takes the Fourier transform of the BOLD signals $y$ at the set of Nyquist frequencies. We set the number of hidden layers to be 3, with hidden widths $M,\lfloor\frac{M}{2}\rfloor,\lfloor\frac{M}{4}\rfloor$. For the marginal likelihood emulator $p_{\gamma}(\cdot|\cdot)$, we use a conditional neural spline normalizing flow (Durkan et al., 2019) with $5$ spline coupling transforms, each with a fully connected 3 layer conditioning network with hidden widths of 64 and ReLU activations. Both networks are trained on synthetic data generated from the simulator outlined in Section 2 using Adam with a batch size of $100$ and learning rate of $10^{-5}$ for $10^{5}$ iterations. Note that separate summary network and likelihood model pairs were trained for the forward models (20) and (21).
@@
For calibrating the simulator hyperparameters using the non-differentiable loss (19), we used Bayesian optimization (Frazier, 2018) under the expected improvement acquisition function for a total of $200$ runs using the rsfMRI from a randomly selected HCP-YA subject and both forward models. The prior smoothness hyperparameters are selected by maximizing (17) over a pre-defined grid. For single subject analysis, this was repeated for each subject. For multi-subject analysis, we use a common set of smoothing parameters for all reconstructions, as is standard practice.
4.4 Competing Hemodynamic Estimators
@@
where $A(\theta)\in\mathbb{R}^{M\times M}$ is a convolution matrix formed from the kernel (20), $s\in\mathbb{R}^{M}$ denotes the discretized neural signals, and $\eta>0$ is a regularization parameter. Optimization is performed using block-coordinate descent, where the (convex) subproblem in $s$ is solved using alternating direction method of multipliers and the (non-convex) subproblem in $\theta$ is solved using golden section search. We set the penalty parameter $\eta=\sigma\sqrt{2\log(M)}$, where $\sigma$ is the calibrated noise level. The second approach, referred to as DUP (deep unrolled prior), adapts Gossard and Weiss (2024) to perform vertex-wise joint reconstruction. Specifically, DUP first pre-trains an unrolled network $\mathcal{A}_{\rho}:\mathbb{R}^{M}\times\mathbb{R}^{M\times M}\mapsto\mathbb{R}^{M}$, parameterized by weights $\rho$, to minimize
@@
via golden section search. We take $\mathcal{A}_{\rho}$ to be a 5-layer unrolled fully connected ReLU network with hidden width $M$ and train on synthetic data from the simulator model in Section 2 using Adam (learning rate $10^{-5}$, $10^{5}$ iterations, batch size $100$). We also attempted to compare our method against Cherkaoui et al. (2021). However, extending their method to a high-resolution surface mesh proved computationally intractable, which is not surprising as the approach was originally developed to work on coarse, atlas-based surface discretizations.
For the two-parameter model (21), we compare our method to Wu et al. (2013), implemented in the MATLAB package rsHRF (Wu et al., 2021). rsHRF uses a two stage procedure, first estimating neural spike times by thresholding the fMRI time series, then estimating hemodynamic parameters conditional on the detected spikes. We used the suggested defaults for all method hyperparameters.
4.5 Evaluation
4.5.1 Hemodynamic Reconstruction
To evaluate the performance of the hemodynamic estimation on synthetic data, we compute the mean squared error (MSE) and bias of the reconstructions, averaged over the mesh. The quality of the uncertainty quantification procedure from Section 3.5 is assessed by the empirical coverage of the $95\%$ pointwise intervals. In the synthetic data experiments, this is computed as the fraction of mesh locations for which the intervals cover the true hemodynamic parameter. For in-vivo data, where we lack access to ground truth, we instead approximate the empirical coverage of the intervals using the HCP test-retest data; where one scan is used to form the intervals and their empirical coverage is calculated over the reconstructions from the other scan. Because hemodynamic response shape is expected to be relatively stable within healthy subjects between multiple scans over a short time period (months), the test-retest scans can be (approximately) considered as independent replications from model (4) with the same $\theta$, though different $s$ and $\epsilon$, and thus can be used to assess the stability of our hemodynamic coupling estimation methodology. We also compute the average interval length over the mesh for use as a scalar measure of reconstruction uncertainty.
4.5.2 Connectivity
Spatial variability in the cortical hemodynamic response can bias downstream functional connectivity estimates (Lindquist et al., 2009; Rangaprakash et al., 2018). We therefore evaluate whether our spatially varying hemodynamic estimates can improve connectivity analysis by comparing seed-based effective connectivity computed from three signal representations: I) raw (non-deconvolved) BOLD signals, II) deconvolved BOLD signals assuming a spatially constant and known hemodynamic response, III) deconvolved BOLD signals under a spatially varying hemodynamic response estimated by our methodology. Under the LTI forward model (3), with a known response kernel (either fixed or pre-estimated), the resulting linear inverse problem for latent signal reconstruction has been widely studied. For simplicity, we choose to perform deconvolution using the Wiener filter (Wiener, 1949), acknowledging that better performance may be obtained through more modern techniques. Using a predefined seed region of interest (ROI), we estimate the effective connectivity from the time series at each non-ROI vertex to the ROI-averaged time series. We use a bivariate Granger Causality (GC) based definition of effective connectivity (Roebroeck et al., 2005) with an autoregressive lag of $2$. For each non-ROI vertex, p-values are calculated from the $F$-tests assessing GC and then corrected to control false discovery rate at the $0.05$ level using the Benjamini-Hochberg procedure (Benjamini and Hochberg, 1995).
4.5.3 Population Analysis
We evaluate inter-subject variation of the hemodynamic field estimates using the functional principal components analysis (fPCA) approach proposed in (Lila et al., 2016). For a cohort of $N$ subjects, fPCA is used to obtain the Karhunen–Loève decomposition of the scalar hemodynamic coupling parameter $\tilde{\theta}_{i}(x_{v})$ for subject $i$ at location $x_{v}$:
5 Results
5.1 Synthetic Data Results
Table 1: Synthetic-data results: estimation performance (MSE, bias) and uncertainty quantification (empirical coverage (EC) and average interval length), averaged across the cortical surface.
@@ 5.1 Synthetic Data Results
Table 1 A) compares the estimation results for the synthetic data using the one-parameter model (20). Our method exhibits mean squared errors and biases that are, on average, an order of magnitude smaller than both JointMAP and DUP. Recall that both of the competitors perform joint reconstruction of signal and hemodynamic parameters, resulting in the estimation of the latter being dependent on the estimation of the former. Since even the non-blind signal reconstruction is ill-posed in this setting, we attribute the poorer performance of the competing methods, at least in part, to the propagation of errors from the signal reconstruction to the target hemodynamic estimation. Our approach avoids this issue by marginalizing the signal out of the data likelihood.
@@
The second key differentiating feature of our method is the incorporation of spatial regularization. To evaluate its effect, we drop the prior from Section 3.2 and apply the learned marginal posterior mean (6), denoted MPM, independently to each mesh BOLD time series. As shown in the third row of Table 1 A), on average, this results in a greater than $6\times$ increase in the MSE compared to our spatially regularized framework. While MPM outperforms both JointMAP and DUP owing to its avoidance of explicit signal reconstruction, we see that incorporating spatial smoothness of the hemodynamics provides substantial additional improvement.
@@
Table 1 B) shows the estimation results for the two-parameter forward model (21). Estimation of $\theta_{1}$ echoes the results from Table 1 A), where our method again results in an order of magnitude reduction in average MSE compared to rsHRF. Although our method exhibits lower MSE for $\theta_{2}$ as well, both approaches perform relatively poorly for this parameter.
@@
Table 1 C) shows the $95\%$ empirical pointwise coverage and average interval length for the procedure from Section 3.5 for both synthetic datasets. For the one-parameter model (20), we see reasonably good empirical coverage, with slight overconfidence. For the two-parameter model (21), the coverage on $\theta_{1}$ indicates a bit of conservatism, though the intervals are quite tight, indicating strong statistical identifiability of this parameter. While the empirical coverage for $\theta_{2}$ is reasonable, the estimation uncertainty is excessively large. Indeed, given $\theta_{2}\in[-1,1]$, the average interval length occupies approximately $70\%$ of the possible values of the parameter. Coupled with the relatively poor estimation performance for both methods for $\theta_{2}$, these results imply there are likely fundamental limitations on its statistical identifiability in resting-state data under HCP-style acquisition and noise level.
5.2 In-vivo Data Results
@@
In this section, we apply our methodology under the one-parameter model (20) to the real resting-state HCP data. We choose this model over model (21) due to the superior statistical identifiability of model parameters reported in Section 5.1.
5.2.1 Hemodynamic Reproducibility
Figure 3: (Left) Estimates of the $\theta$-field for the one-parameter hemodynamic coupling model (20) on the HCP test-retest data for three randomly selected subjects. Inter-subject variability is noticeably larger than within-subject between-scan variability. (Right) Length of the $95\%$ pointwise intervals calculated from the procedure in Section 3.5 on the Scan 1 data.
@@ 5.2.1 Hemodynamic Reproducibility
The spatial maps of $\widehat{\boldsymbol{\theta}}$ for 3 randomly selected HCP test-retest subjects are shown in the left part of Figure 3. Qualitatively, we see that between-subject variability is noticeably larger than within-subject (between-scan) variability for each subject, with many unique subject-specific spatial features preserved across scans. This result is also apparent quantitatively, as the average within-subject squared norm of the difference between scan 1 and scan 2 ($\approx 0.0214$) was substantially lower than the average between-subject squared norm ($\approx 0.0590$).
@@
For each subject in Figure 3, we ran the uncertainty quantification procedure from Section 3.5 on the scan 1 data to construct pointwise $95\%$ intervals for the reconstructed field. The resulting interval lengths are shown in the right part of Figure 3. We notice substantial spatial variability in these maps, indicating the reconstruction uncertainty is highly location dependent. We evaluated the empirical coverage of these intervals on the estimates from the scan 2 data. The average empirical coverage was $\approx 98.5\%$, indicating good calibration with mild conservatism.
5.2.2 Population Analysis